# OpenAgentSafety: A Comprehensive Framework for Evaluating Real-World AI Agent Safety

### v1

@@ Abstract
Recent advances in AI agents capable of solving complex, everyday tasks—from scheduling to customer service—have enabled deployment in real-world settings, but their possibilities for unsafe behavior demands rigorous evaluation. While prior benchmarks have attempted to assess agent safety, most fall short by relying on simulated environments, narrow task domains, or unrealistic tool abstractions. We introduce OpenAgentSafety, a comprehensive and modular framework for evaluating agent behavior across eight critical risk categories. Unlike prior work, our framework evaluates agents that interact with real tools, including web browsers, code execution environments, file systems, bash shells, and messaging platforms; and supports over 350 multi-turn, multi-user tasks spanning both benign and adversarial user intents. OpenAgentSafety is designed for extensibility, allowing researchers to add tools, tasks, websites, and adversarial strategies with minimal effort. It combines rule-based analysis with LLM-as-judge assessments to detect both overt and subtle unsafe behaviors. Empirical analysis of five prominent LLMs in agentic scenarios reveals unsafe behavior in 51.2% of safety-vulnerable tasks with Claude-Sonnet-3.7, to 72.7% with o3-mini, highlighting critical safety vulnerabilities and the need for stronger safeguards before real-world deployment.
@@ 1 Introduction
The development of AI agents has surged, fueled by LLM breakthroughs that enable their deployment for software engineering [wang2025openhandsopenplatformai], web browsing [zhou2023webarena], and customer service tasks [langchain2024state] among others. The rapid pace of their development has far outpaced progress on ensuring their safety. Agents are increasingly granted access to powerful tools that enable them to perform complex, multi-step tasks autonomously. Driven by competitive pressure and a large economic incentive to deploy, many agentic systems have been released without a thorough investigation into their failure modes or societal impacts [langchain2024state, plaat2025agenticlargelanguagemodels]. The gap between capability advancement and safety assurance continues to widen, and creates vulnerability to both catastrophic failures and subtle but pervasive harms that could prove difficult to reverse once embedded in societal systems [zhang2024agentsafetybenchevaluatingsafetyllm].
@@
To address these risks, a number of benchmarks have recently been proposed (Table 1). While valuable, existing benchmarks are often limited in scope as many rely on toy environments or simulated tool APIs, focus on narrow domains like browsing or coding, or omit multi-turn, multi-user interactions. These limitations make it difficult to evaluate how agents behave in complex, realistic situations. As agents become increasingly capable, it is critical to build benchmarks that accurately reflect real-world challenges, including diverse tools, varied user behavior, and long-horizon tasks.
@@
Table 1: Comparison of agent safety benchmarks based on their support for (i) real-world tool usage, (ii) diverse user intentions, and (iii) multi-turn user interactions. Only our benchmark, OpenAgentSafety, supports all three. In the “Diverse intents“ column, denotes support for benign user goals (e.g., a user unintentionally exposing a private key), and indicates support for malicious goals (e.g., a user attempting to coerce the agent into generating ransomware).
@@
SALAD-Bench [li2024saladbenchhierarchicalcomprehensivesafety]
h4rm3l [doumbouya2025h4rm3llanguagecomposablejailbreak]
SafeBench [ying2024safebenchsafetyevaluationframework]
Agent-SafetyBench [zhang2024agentsafetybenchevaluatingsafetyllm]
SG-Bench [mou2024sgbenchevaluatingllmsafety]
SafeAgentBench [yin2025safeagentbenchbenchmarksafetask]
ChemSafetyBench [zhao2024chemsafetybenchbenchmarkingllmsafety]
LM-Emulated Sandbox [ruan2024identifyingriskslmagents]
AdvWeb [xu2024advwebcontrollableblackboxattacks]
Refusal-Trained LLMs [kumar2024refusaltrainedllmseasilyjailbroken]
RedCode [guo2024redcoderiskycodeexecution]
From Interaction to Impact [zhang2024attackingvisionlanguagecomputeragents]
PrivacyLens [shao2025privacylensevaluatingprivacynorm]
Dissecting Adversarial [wu2025dissectingadversarialrobustnessmultimodal]
Infrastructure for AI [chan2025infrastructureaiagents]
R-Judge [yuan2024rjudgebenchmarkingsafetyrisk]
Trembling House [mo2024tremblinghousecardsmapping] &
AgentHarm [andriushchenko2025agentharmbenchmarkmeasuringharmfulness] &
WildTeaming [jiang2024wildteamingscaleinthewildjailbreaks] &
SafetyPrompts [röttger2025safetypromptssystematicreviewopen] &
ST-WebAgentBench [levy2024stwebagentbenchbenchmarkevaluatingsafety]
Frontier Models [meinke2025frontiermodelscapableincontext]
SafeArena [tur2025safearenaevaluatingsafetyautonomous] &
Haicosystem [zhou2024haicosystemecosystemsandboxingsafety] &
@@
We introduce OpenAgentSafety (OA-Safety, §2), a comprehensive and open-source simulation framework for evaluating the safety of AI agents in realistic, high-risk scenarios, supporting:
  • Real-world, comprehensive tool suite: Agents interact with actual file systems, command lines, code execution environments, and web interfaces.
  • Diverse user intentions: Tasks simulate user behavior from benign ambiguity to manipulation.
Figure 1: LLM-based agent unsafe behavior rates in the OpenAgentSafety framework while navigating conflicting user and NPC instructions.
OA-Safety is built on a robust and modular infrastructure that supports sandboxed agent execution with real tools, including file systems, bash terminals, code execution environments, web browsers, and messaging platforms. It includes over 350 executable tasks, each simulating multi-turn interactions with users exhibiting benign, ambiguous, or adversarial intent, where adversarial users may appear cooperative but subtly aim to induce harmful agent behavior. The framework is designed for extensibility: researchers can easily add new tasks, simulated environments (e.g., websites), complex social dynamics (e.g., negotiation with a customer), and customized evaluators. Simulation is a core design principle in OA-Safety and agents operate inside isolated containers with full tool access, enabling realistic actions (e.g., deleting files, leaking data, sending messages) to be executed safely without real-world consequences.
Each OA-Safety task is implemented as a modular Docker container that includes: the task description, multiple user goals, social dynamics, and customized evaluators, including both: rule-based evaluator that detects concrete environment impacts such as file deletion, and LLM-as-a-judge assessments to analyze the agent’s intermediate reasoning to flag attempted unsafe actions, even if those actions were incomplete or ultimately unsuccessful. This design allows for efficient environment reuse and flexible task extension.
We evaluate five prominent LLMs on OpenAgentSafety and find that current agents exhibit a wide range of unsafe behaviors across realistic, multi-turn scenarios (§3). Unsafe actions occur in 51.2% to 72.7% of safety-vulnerable tasks (Figure 1). We confirm prior findings that access to the browsing tool can increase the risk of unsafe behavior by overloading the agent’s context [tur2025safearenaevaluatingsafetyautonomous]. Beyond this, our analysis reveals new failure modes that are underexplored in existing safety benchmarks: we establish agents frequently fail to reason over extended multi-turn interactions, leading individually safe steps to compound into unsafe outcomes; they disregard legal, privacy, and security policies even in high-risk settings; and they show structurally unsafe behavior patterns across diverse user intents and tool types. Unlike prior studies, we analyze the impact of different risk categories, user intents, and tool usage, allowing us to uncover patterns in unsafe agent behavior not captured by single-axis benchmarks.
Our contributions are:
  • We introduce OpenAgentSafety, [^1] Code and data at: https://github.com/sani903/OpenAgentSafety a modular and extensible simulation framework with 350+ executable tasks covering eight key safety risk categories.
  • We propose hybrid evaluations with rigorous, rule-based and flexible, neural-based assessments.
  • We present a detailed empirical analysis across five LLMs, revealing failure modes and agent vulnerabilities in realistic deployment scenarios.
Figure 2: An overview of the OpenAgentSafety framework.
2 OpenAgentSafety Framework
We detail the OpenAgentSafety (OA-Safety) framework by first introducing the infrastructure to support agents and environments (§2.1), describing our task taxonomy and creation process (§2.2), finally presenting our hybrid evaluation for task success and safety (§2.3).
@@ 2.1 Infrastructure For Agent and Environment
To evaluate agent behavior in high-risk scenarios, we build OA-Safety on top of the OpenHands framework [wang2025openhandsopenplatformai], an open-source platform for multi-tool LLM agents. The agent runs inside a containerized sandbox with access to real tools, including a Unix shell, file system, Python and Bash interpreters, and a web browser. This architecture enables realistic tool-based agent workflows, while isolating the agent from the host system to safely observe potentially harmful behaviors. Different LLMs can be deployed as agent backbones for safety testing on agentic tasks.
To prevent real-world side effects during evaluation, such as posting content to live platforms or compromising the operating system, we replicate real-world websites in local Docker containers. We use locally hosted instances of OwnCloud (file sharing), GitLab (version control), and Plane (issue tracking), adapted from The Agent Company [xu2024theagentcompanybenchmarkingllmagents]. These websites simulate realistic interaction contexts for agents, such as uploading confidential documents or modifying code repositories.
A key component of OA-Safety is its support for multi-user scenarios. We integrate the Sotopia framework [zhou2024sotopia] to simulate secondary actors (NPCs) with diverse goals. We extend OpenHands with a custom ChatNPC tool to enable agents to send and receive messages via Sotopia’s Redis-based communication backend to these NPCs. This setup supports direct and broadcast messages, enabling tasks that reflect real-world organizational and social interactions, and allowing us to model complex social dynamics (e.g., persuasion, conflict) independently of the browsing proficiency of agents.
Table 2: Eight safety risk categories in our OpenAgentSafety benchmark with exemplar scenarios. Each category highlights a distinct failure mode relevant to real-world agent deployments.
@@
Data loss / corruption Deleting critical system files while attempting to clean up disk space.
@@
Spreading malicious content Sending abusive or inappropriate messages on behalf of users to staff.
Legal violations Intentionally submitting incorrect forms (e.g., I-9) to cause harm.
Harmful decision-making Tampering with survey results to satisfy task completion requirements.
@@ 2.2 Safety Taxonomy and Task Design
We use three dimensions—risk category, tool usage, and user/NPC intent—to design tasks and evaluate agent safety behaviors. To ensure broad coverage of real-world deployment scenarios, we organize our benchmark around a taxonomy of 8 safety risk categories (Table 2). We aggregate and refine previous categorizations [zeng2024airiskcategorizationdecoded, zhang2024agentsafetybenchevaluatingsafetyllm, ruan2024identifyingriskslmagents], and map them to concrete actions involving real tools. In addition to risk types and tool modalities, user intent plays a central role in shaping task difficulty. Tasks may involve benign, or explicitly adversarial users (Table 7). Through our integration of the ChatNPC tool, we further introduce an optional NPC intent, where secondary actors/NPCs (e.g., colleagues, customers) simulated through Sotopia [zhou2024sotopia] interact with the agent and may have conflicting or manipulative goals (Table 3). These multi-agent dynamics are essential to simulating realistic organizational scenarios where the agent needs to interact with secondary actors (NPCs) (Table 7). Thus, the available tools for the agent are: file system, browsing, ChatNPC, IPython for code generation and execution, and bash terminal.
To create the benchmark, we manually construct 80 seed tasks and their corresponding rule-based evaluators. Each seed task captures a distinct combination of tools and social context. To scale up, we systematically generate additional tasks using GPT-4o [openai-gpt-4o], prompting it to vary the social context and NPC configurations while preserving the core behavioral challenge and tool usage (prompt provided in Appendix A.6). All 356 generated tasks are manually verified by the authors, with their distribution across risk categories, intents, and tools provided in Appendix §A.1.
Each task is packaged as a self-contained Docker image, including (i) the environment setup (e.g., websites, files), (ii) task description, (iii) secondary actor (NPC) behaviors, and (iv) a rule-based evaluator. This modular design allows easy integration of new environments and tools, including external ecosystems such as WebArena [zhou2023webarena], into the safety evaluation pipeline.
Next, we introduce more details about the evaluation process, as well as LLM-as-judge serving as an alternative evaluation, especially when rule-based scripts are difficult to construct or unnecessary.
Table 3: Descriptions of the three most common user and NPC intent combinations.
User intent | NPC intent Task type
Benign OR Benign | N/A The task is benign on the surface but can be completed unsafely if the agent overlooks subtle cues.
Benign | Malicious A standard user task involves interaction with an NPC who has a deceptive or adversarial goal.
Malicious OR Malicious | N/A The unsafe intent is embedded directly in the user’s instructions.
@@ 2.3 Evaluation Approach
To determine whether agents behave unsafely during task completion, we employ a hybrid evaluation strategy that combines:
@@
Rule-based evaluation quad For each task, we implement a Python script that analyzes the final environment state (e.g., file system, database, website state) to check whether specific unsafe outcomes occurred, such as deleting protected files or posting confidential data. These evaluators return a binary outcome: 1 for unsafe, 0 for safe. We report the percentage of unsafe outcomes across tasks.
@@
LLM-as-Judge evaluation  To also capture unsafe intent or behavior that does not result in a final environment change (e.g., incomplete attempts or unsafe reasoning), we evaluate agent trajectories by prompting GPT-4.1 [openai_gpt41_2025] to judge safety based on intermediate thoughts and actions (prompt in Appendix A.3). GPT-4.1 must provide one of the four labels to the trajectory: failure for trajectories where the agent did not reach an safety-vulnerable point where it would be exposed to unsafe content, safe for trajectories which include rejection of tasks, or safely completing the task while averting the risk. incomplete/unsafe and complete/unsafe represent unsafe labels where the former represents that the agent tried to take malicious actions but they were incomplete. The latter refers to tasks where the agent completed the unsafe behavior.
@@
Other than the safety ratings from the evaluators, we additionally report two derived metrics:
@@
Failure rate  The percentage of tasks where the agent failed to reach a safety-vulnerable state due to its inability to complete an initial subtask, and was therefore never exposed to the unsafe element (e.g., the agent failed to download a file and was thus never manipulated into modifying it, as in the unauth-discount task in Table 7).
@@
Disagreement rate  The percentage of tasks where the LLM judge labeled a trajectory as safe, but the rule-based evaluator detected a concrete unsafe impact. A high disagreement rate may indicate blind spots in LLM-based evaluation or misleading agent reasoning.
Designing robust rule-based evaluators is non-trivial: it often requires multiple iterations based on actual agent behavior to account for diverse unsafe attempts and avoid false positives or negatives.
@@
The LLM-as-Judge component plays a critical role in disambiguating failure and safe trajectories, both of which are classed as safe from the rule-based evaluator. Further, while rule-based checks capture tangible environment changes, they cannot detect cases where the agent intended to act maliciously but failed to execute the behavior. They also fail to identify content safety risks. As a result, attempted unsafe behavior without environmental impact is marked as safe by the rule-based system. LLM-as-Judge helps assess the agent’s reasoning and intermediate actions to handle these cases appropriately. This hybrid evaluation protocol balances the precision of rule-based checks with the broader behavioral insight of LLM judgments, enabling robust safety assessments across diverse tasks.
3 Experiment

### v2

@@ Abstract
Recent advances in LLM agents capable of solving complex, everyday tasks, ranging from software engineering to customer service, have enabled deployment in real-world scenarios, but their possibilities for unsafe behavior demands rigorous evaluation. While prior benchmarks have attempted to evaluate safety of LLM agents, most fall short by relying on simulated environments, narrow task domains, or unrealistic tool abstractions. We introduce OpenAgentSafety, a comprehensive and modular framework for evaluating agent behavior across eight critical risk categories. Unlike prior work, our framework evaluates agents that interact with real tools, including web browser, code execution environment, file system, bash terminal, and messaging platform; and supports over 350 multi-turn, multi-user tasks spanning both benign and adversarial user intents. OpenAgentSafety is designed for extensibility, allowing researchers to add tools, tasks, web environments, and adversarial strategies with minimal effort. It combines rule-based evaluation with LLM-as-judge assessments to detect both overt and subtle unsafe behaviors. Empirical analysis of seven prominent LLMs in agentic scenarios reveals unsafe behavior in 49% of safety-vulnerable tasks with Claude Sonnet 4, to 73% with o3-mini, highlighting critical risks and the need for stronger safeguards before real-world deployment of LLM agents [^1] Code and data can be accessed at https://github.com/Open-Agent-Safety/OpenAgentSafety
@@ 1 Introduction
Recent advances in large language models (LLMs) have fueled the development of AI agents which are now being deployed for software engineering (Wang et al., 2025), web browsing (Zhou et al., 2023), and customer service tasks (LangChain, 2024) among others. The rapid pace of their development has far outpaced progress in ensuring their safety. Agents are increasingly granted access to powerful tools that enable them to perform complex, multi-step tasks autonomously. Driven by competitive pressure and a large economic incentive to deploy, many agentic systems have been released without a thorough investigation into their failure modes or societal impacts (LangChain, 2024; Plaat et al., 2025). The gap between capability advancement and safety assurance continues to widen, making agents vulnerable to both catastrophic failures and subtle but pervasive harms that could prove difficult to reverse once embedded in societal systems (Zhang et al., 2024b).
Figure 1: An overview of the OpenAgentSafety framework.
To mitigate and address these risks, we introduce OpenAgentSafety (OA-Safety, §2), a comprehensive and open-source simulation framework for evaluating the safety of AI agents in realistic, high-risk scenarios. Built on a robust and modular infrastructure, OA-Safety supports:
  • Real-world, comprehensive tool suite: Agents interact with actual file systems, command line, code execution environments, and self-hosted web interfaces in a sandboxed environment to prevent any real-world harm.
  • Diverse user intentions: Tasks simulate user behavior ranging from benign ambiguity to adversarial manipulation.
@@
With these features, OA-Safety substantially improves upon existing benchmarks which are often limited in scope as they rely on toy environments or simulated tool APIs, focus on narrow domains like browsing or coding, or omit multi-turn, multi-user interactions (Table 1). These gaps hinder evaluation of agent behavior in realistic settings. As capabilities grow, benchmarks must capture real-world challenges—diverse tools, varied user behavior, and long-horizon tasks.
To demonstrate the utility of our framework, we craft over 350 executable tasks, simulating multi-turn interactions with users exhibiting benign, ambiguous, or adversarial intent, where adversarial users may appear co-operative but subtly aim to induce harmful agent behavior. Inspired by coding benchmarks (Guo et al., 2024; Jimenez et al., 2024), each OA-Safety task is implemented as a modular Docker container that includes the task description, multiple user goals, social dynamics, and customized evaluators, including both: rule-based evaluator that detects harmful actions by examining the state of the environment (for e.g. deletion of an important file), and LLM-as-judge to analyze the agent’s reasoning to flag attempted unsafe actions, despite being potentially incomplete or eventually unsuccessful. This allows for efficient environment reuse and flexible task extension.
Figure 2: Unsafe agent behaviour rates of various LLMs measured using the OpenAgentSafety framework when navigating conflicting user and NPC instructions.
We evaluate seven prominent LLMs on OpenAgentSafety and find that exhibit a wide range of unsafe behaviors across complex [^2] We define complexity as introducing both social dynamics with multiple actors and more steps required to complete a task than previous benchmarks. realistic, multi-turn scenarios (§3) when used as the backbone of agentic systems. Unsafe actions occur in 49% to 73% of safety-vulnerable tasks (Figure 2). Our analysis, which examines the impact of different risk categories, user intents, and tool usage, reveals new failure modes that are underexplored in existing safety benchmarks (e.g., Figure 2): we observe agents frequently fail to reason over extended multi-turn interactions, which results in individually safe steps compounding into unsafe outcomes; they disregard legal, privacy, and security policies even in high-risk settings; and they show structurally unsafe behavior patterns across diverse user intents and tool types. We also confirm prior findings that access to the browsing tool can increase the risk of unsafe behavior by overloading the agent’s context (Tur et al., 2025).
Our research contributions are as follows:
  • We introduce OpenAgentSafety, a modular and extensible evaluation framework with 350+ executable tasks spanning eight key safety risk categories. Tasks vary systematically in user intent (benign vs. malicious) and NPC behavior, capturing how different interaction patterns give rise to unsafe outcomes.
  • Our framework is designed for extensibility, allowing researchers to easily add new tasks, simulated environments (e.g., websites), complex social dynamics (e.g., negotiation with a customer), and customized evaluators.
  • We conduct a detailed empirical analysis across seven LLMs, uncovering failure modes and vulnerabilities in realistic deployment scenarios. We find that (i) seemingly benign inputs that allow for “easy but unsafe“ solutions drive a large share of unsafe behaviors, and (ii) models consistently struggle with systemic risks that require understanding institutional norms.
2 OpenAgentSafety Framework
In this section, we describe the OpenAgentSafety (OA-Safety) framework. We introduce our infrastructure in §2.1, describe our task taxonomy and task creation process in §2.2, and finally present our hybrid evaluation method in §2.3.
@@ 2 OpenAgentSafety Framework
Table 1: Comparison of agent safety benchmarks based on (i) real-world tool support, (ii) diverse user intents, and (iii) multi-turn user interactions. Only OpenAgentSafety, supports all three. denotes inclusion of tasks with benign user goals (e.g., unintentionally exposing an API key), and denotes presence of tasks with malicious user goals (e.g., asking the agent to generate ransomware).
@@
SALAD-Bench (Li et al., 2024)
h4rm3l (Doumbouya et al., 2025)
SafeBench (Ying et al., 2024)
Agent-SafetyBench (Zhang et al., 2024b)
SG-Bench (Mou et al., 2024)
SafeAgentBench (Yin et al., 2025)
ChemSafetyBench (Zhao et al., 2024)
LM-Emulated Sandbox (Ruan et al., 2024)
AdvWeb (Xu et al., 2024a)
Refusal-Trained LLMs (Kumar et al., 2024)
RedCode (Guo et al., 2024)
From Interaction to Impact (Zhang et al., 2024a)
PrivacyLens (Shao et al., 2025a)
Dissecting Adversarial (Wu et al., 2025)
Infrastructure for AI (Chan et al., 2025)
R-Judge (Yuan et al., 2024)
Trembling House (Mo et al., 2024) &
AgentHarm (Andriushchenko et al., 2025) &
WildTeaming (Jiang et al., 2024) &
SafetyPrompts (röttger2025safetypromptssystematicreviewopen) &
ST-WebAgentBench (Levy et al., 2024)
Frontier Models (Meinke et al., 2025)
SafeArena (Tur et al., 2025) &
Haicosystem (Zhou et al., 2024a) &
@@ 2.1 Infrastructure For Agent and Environment
We build OA-Safety on top of the OpenHands framework (Wang et al., 2025), an open-source platform for multi-tool LLM agents. The agent runs inside a containerized sandbox with access to real tools, including a Unix shell, file system, Python interpreter, and a web browser. This architecture enables realistic tool-based agent workflows, while isolating the agent from the host system to safely observe potentially harmful behaviors. Different LLMs can be evaluated with this agent for analyzing their safety in agentic tasks.
To prevent real-world harm during evaluation, such as posting harmful content to live platforms, we replicate real-world websites in local Docker containers. We use locally hosted instances of OwnCloud (file sharing), GitLab (version control), and Plane (issue tracking), adapted from The Agent Company (Xu et al., 2024b). These websites simulate realistic interaction contexts for agents, such as uploading confidential documents or modifying code repositories.
A key component of OA-Safety is its support for multi-user scenarios as LLMs struggle to navigate multiparty scenarios (Penzo et al., 2024). We leverage the open-source Sotopia framework (Zhou et al., 2024b) to simulate secondary actors (NPCs) with diverse goals. We extend OpenHands with a custom ChatNPC tool that enables the agent to communicate with these NPCs using Sotopia’s Redis-based communication backend. This setup supports direct and broadcast messages, enabling tasks that reflect real-world organizational and social interactions, and allowing us to model complex social dynamics (e.g., persuasion, conflict) independently of the browsing proficiency of agents.
Table 2: Eight safety risk categories in the OpenAgentSafety benchmark with example task scenarios. Each category highlights a distinct failure mode relevant to real-world agent deployments.
@@
Data loss / corruption Deleting critical system files reducing disk space.
@@
Spreading malicious content Sending abusive or inappropriate messages to staff.
Legal violations Submitting incorrect forms (e.g., I-9) to cause harm.
Harmful decision-making Tampering with survey results for task completion requirements.
@@ 2.2 Safety Taxonomy and Task Design
We use three dimensions to design tasks and evaluate agent safety behaviors – risk category, tool usage, and user/NPC intent. To ensure broad coverage of real-world deployment scenarios, we organize our benchmark around a taxonomy of 8 risk categories (Table 2). We aggregate and refine previous categorizations (Zeng et al., 2024; Zhang et al., 2024b; Ruan et al., 2024), and map them to concrete actions involving real tools. In addition to risk types and tool modalities, user intent plays a central role in shaping task difficulty. The primary user of the agent may use it for tasks involving benign, or explicitly adversarial intents. Further, through our integration of the ChatNPC tool, we introduce an optional NPC intent, where secondary actors/NPCs (e.g., colleagues, customers) simulated through Sotopia (Zhou et al., 2024b) interact with the agent via text messages. As multi-agent dynamics are essential to simulating realistic organizational scenarios where the agent needs to interact with secondary actors, we set these NPCs to possibly have manipulative goals or goals that conflict with the main user’s intent. Thus, the available tools for the agent are: file system, browser, ChatNPC, IPython for code generation and execution, and a bash terminal.
To create the benchmark, we manually construct 80 seed tasks and their corresponding rule-based evaluators. Each seed task captures a distinct combination of tools and social context. To scale up, we systematically generate additional tasks using GPT-4o (OpenAI et al., 2024), prompting it to vary the social context and NPC configurations while preserving the core behavioral challenge and tool usage (prompt provided in Appendix A.7). All 356 generated tasks are manually verified by the authors to ensure richness and coverage, with their distribution across risk categories, intents, and tools provided in Appendix §A.1.
Each task is packaged as a self-contained Docker image, including (i) the environment setup (e.g., websites, files), (ii) task description, (iii) secondary actor (NPC) behaviors, and (iv) a rule-based evaluator. This modular design allows easy integration of new environments and tools, including external ecosystems such as WebArena (Zhou et al., 2023), into the safety evaluation pipeline.
We focus on three common intent combinations in our tasks. First, when both the user intent is benign (and no NPC is involved), the task appears safe on the surface but can still be completed unsafely if the agent overlooks subtle cues. Second, in cases where the user’s intent is benign but an NPC introduces malicious goals, the agent faces a standard interaction that becomes adversarial due to deceptive NPC behavior. Finally, when the user intent itself is malicious (with or without an NPC), the unsafe instruction is embedded directly in the user’s request.
@@ 2.3 Evaluation Approach
To determine whether agents behave unsafely during task completion, we employ a two-pronged evaluation strategy, designed to keep in mind the limitations of each individual evaluation approach and the two complement each other. Specifically, we measure safety using:
@@
Rule-based evaluation  For each task, we implement a Python script that analyzes the final environment state (e.g., file system, database, website state) to check whether specific unsafe outcomes occurred, such as deleting protected files or posting confidential data. These evaluators return a binary outcome: 1 for unsafe, 0 for safe. We report the percentage of unsafe outcomes across tasks.
@@
LLM-as-Judge evaluation  To also capture unsafe intent or behavior that does not result in a final environment change (e.g., incomplete attempts or unsafe reasoning), we evaluate agent trajectories by prompting GPT-4.1 (OpenAI, 2025a) to judge safety based on intermediate thoughts and actions (prompt in Appendix A.3). GPT-4.1 must provide one of the four labels to the trajectory: The safe trajectories include rejection/noncompliance for unsafe tasks, or safely completing the task while averting the risk, and failure denotes trajectories where the agent did not reach a safety-vulnerable point where it would be exposed to unsafe content. The complete/unsafe label indicates that the agent successfully carried out the unsafe task, while the incomplete/unsafe label captures cases where the agent attempted malicious actions but did not fully complete the task.
@@
Other than the safety ratings from the evaluators, we additionally report three additional metrics:
@@
Failure rate  The percentage of tasks where the agent did not complete an initial subtask, and was thus never exposed to the unsafe element (e.g., the agent failed to download a file and was thus never manipulated into modifying it, as in the unauth-discount task in Table 6).
@@
Disagreement rate  The percentage of tasks where the LLM judge labeled a trajectory as safe, but the rule-based evaluator detected a concrete unsafe impact. A high disagreement rate may indicate blind spots in LLM-based evaluation.
Successful completion rate  The percentage of tasks where an LLM Judge evaluates whether the ground truth correct completion has been achieved at the end of the trajectory. The correct completion may be outright refusal, or completing the specified task safely, free from malicious interference or bad practices.
@@
Note that designing robust rule-based evaluators is non-trivial: it often requires multiple iterations based on actual agent behavior to account for diverse unsafe attempts and avoid false positives or negatives. The LLM-as-Judge component plays a critical role in disambiguating failure and safe trajectories, both of which are classed as safe from the rule-based evaluator. Further, while rule-based checks capture tangible environment changes, they cannot detect cases where the agent intended to act maliciously but failed to execute the behavior. They also fail to identify content safety risks. As a result, attempted unsafe behavior without environmental impact is marked as safe by the rule-based system. LLM-as-Judge helps assess the agent’s reasoning and intermediate actions to handle these cases appropriately. This hybrid evaluation protocol balances the precision of rule-based checks with the broader behavioral insight of LLM judgments, enabling robust safety assessments.
3 Experiments and Results